OpenStudy (anonymous):

Solve: 7 – 5(2 – x) = –(x – 8)

OpenStudy (anonymous):

Distribute the -5= 7-10+5x Distribute the -1 -x+8 So... 7-10+5x=-x+8 Can you solve it from here?

OpenStudy (anonymous):

yes thanks

OpenStudy (anonymous):

Move the variables to one side of the equation, and the numbers to the other side. 5x+x=8-7+10 Simplify 6x=11 6x/6=11/6 x=11/6